Publication List

佐々木建昭（Tateaki Sasaki）

[1] 著作

(1.9) 佐々木建昭ほか、 “新版情報処理ハンドブック”、情報処理学会編、オーム社、東京、 1995。

(1.8) 佐々木建昭・今井浩・浅野孝夫・杉原厚吉、 “計算代数と計算幾何（岩波講座応用数学 [方法９]）”、岩波書店、東京、 1993。

(1.7) 佐々木建昭ほか、 “現代数理科学事典”、大阪書籍、大阪、 1991。

(1.6) 佐々木建昭ほか、 “岩波情報科学辞典”、岩波書店、東京、 1990。

(1.5) 佐々木建昭ほか、 “AI 事典”、 UPU、東京、 1988。

(1.4) 佐々木建昭ほか、 “R&D コンピューティング要覧”、サイエンスフォーラム、東京、 1988。

(1.3) 佐々木建昭・元吉文男・渡辺隼郎、 “数式処理システム（ソフトウェア講座３６巻）”、昭晃堂、東京、 1986。

(1.2) 一松信・山本純恭・内野正弘・佐々木建昭・長尾真・田丸啓吉、 “数と式と文の処理（岩波講座情報科学２３巻）”、岩波書店、東京、 1981。

(1.1) 佐々木建昭、 “数式処理（情報処理草書７巻）”、情報処理学会編、オーム社、東京、 1981。

[2] 近似的代数計算に関する論文

(2.12) T. Sasaki & F. Kako, “Solving multivariate algebraic equation by Hensel construction”, Japan J. Indus. Appl. Math. (to appear).

(2.11) T. Sasaki & T. Yamaguchi, “An analysis of cancellation error in multivariate Hensel construction with floating-point number arithmetic”, Proc. ACM Symposium on Symbolic and Algebraic Computation, ACM (New York), 1998, pp. 1-8.

(2.10) Y. Ozaki & T. Sasaki, “Univariate factor separation and its application to multiple/close root problem”, 数式処理 (日本数式処理学会)、第６巻、 pp. 30-46 (1998).

(2.9) T. Sasaki & M. Sasaki, “Polynomial remainder sequence and approximate GCD”, ACM SIGSAM Bulletin, Vol. 31, pp. 4-10 (1997).

(2.8) K. Shiihara and T. Sasaki, “Analytic continuation and determination of Riemann surface of algebraic functions by computer”, Japan J. Indus. Appl. Math., Vol. 15, pp. 107-116 (1996).

(2.7) T. Sasaki, “A study of approximate polynomials, I – representation and arithmetic –”, Japan J. Indus. Appl. Math., Vol. 12, pp. 137-161 (1995).

(2.6) T. Sasaki, T. Saito and T. Hilano, “Analysis of approximate factorization algorithm I”, Japan J. Indus. Appl. Math., Vol. 9, pp. 351-368 (1992).

(2.5) T. Sasaki, M. Suzuki, M. Kolar and M. Sasaki, “Approximate factorization of multivariate polynomials and absolute irreducibility testing”, Japan J. Indus. Appl. Math., Vol. 8, pp. 357-375 (1991).

(2.4) M. Noda and T. Sasaki, “Approximate GCD and its application to ill-conditioned algebraic equations”, J. Comput. Appl. Math., Vol. 38, pp. 335-351 (1991).

(2.3) M. Ochi, M. Noda and T. Sasaki, “Approximate greatest common divisor of multivariate polynomials and its application to ill-conditioned system of algebraic equations”, J. Inf. Proces., Vol. 14, pp. 292-300 (1991).

(2.2) T. Sasaki and M. Sasaki, “Analysis of accuracy decreasing in polynomial remainder sequence with floating-point number coefficients”, J. Inf. Proces., Vol. 12, pp. 394-403 (1989).

(2.1) T. Sasaki and M. Noda, “Approximate square-free decomposition and root-finding of ill-conditioned algebraic equations”, J. Inf. Proces., Vol. 12, pp. 159-168 (1989).

[3] 数式処理／計算代数に関する論文

(3.27) T. Sado, T. Halaikawa & T. Sasaki, “Pseudo-parallel execution of modular computation”, ACM SIGSAM Bulletin, Vol. 30, No. 1, pp. 4-8.

(3.26) T. Sasaki and M. Sasaki, “A unified method for multivariate polynomial factorizations”, Japan J. Indus. Appl. Math., Vol. 10, pp. 21-39 (1993).

(3.25) M. Kolar and T. Sasaki, “Multivariate quotient by power-series division”, ACM SIGSAM Bulletin, Vol. 26, No. 3, pp. 17-20 (1992).

(3.24) T. Sasaki and M. Sasaki, “On integer-to-rational conversion algorithm”, ACM SIGSAM Bulletin, Vol. 26, No. 2, pp. 19-21 (1992).

(3.23) T. Sasaki and M. Suzuki, “Three new algorithms for multivariate polynomial GCD”, J. Symb. Comp., Vol. 13, pp. 395-411 (1992).

(3.22) 三枝義典、阿部昭博、佐々木建昭、増永良文、元吉文男、佐々木睦子、 “数式処理システム GAL における数学公式データベースのインデキシング手法”、電子情報通信学会論文誌、第 J74-D-I 巻、 pp. 577-585 (1991).

(3.21) T. Sasaki, Y. Fukui, M. Suzuki and M. Sato, “Proposal of a scheme for linking different computer languages – from the viewpoint of algebraic-numeric computation –”, J. Inf. Proces., Vol. 13, pp. 529-533 (1990).

(3.20) 佐々木建昭、竹島卓、 “グレブナ基底の並列計算と連立代数方程式”、情報処理学会論文誌、第 30 巻, pp. 1555-1561 (1989).

(3.19) T. Sasaki and T. Takeshima, “A modular method for Gröber basis construction over Q and solving a system of algebraic equations”, J. Inf. Proces., Vol. 12, pp. 371-379 (1989).

(3.18) T. Sasaki, “Some algebraic algorithms based on head term elimination over polynomial rings”, Lect. Notes in Comp. Sci., Vol. 378, pp. 348-354, 1989.

(3.17) M. Suzuki, T. Sasaki, M. Sato and Y. Fukui, “A hybrid algebraic-numeric system ANS and its preliminary implementation”, Lect. Notes in Comp. Sci., Vol. 378, pp. 163-171 (1989).

(3.16) 佐々木建昭、増永良文、三枝義典、阿部昭博、元吉文男、佐々木睦子、 “国産数式処理システム GAL における数学公式データベース”、第２９回プログラミングシンポジウム報告集、情報処理学会、 1988, pp. 163-173.

(3.15) S. Moritsugu, A. Furukawa, H. Kobayashi and T. Sasaki, “On the power-series solution of a system of algebraic equations”, ACM SIGSAM Bulletin, Vol. 21, pp. 14-23 (1987).

(3.14) H. Kobayashi, A. Furukawa and T. Sasaki, “Gröbner basis of ideals of convergent power series”, Proc. ACM Symposium on Symbolic and Algebraic Computation, ACM (New York), 1986, pp. 225-227.

(3.13) A. Furukawa, T. Sasaki and H. Kobayashi, “Gröbner basis of a module over K{x_{1},…, x_{n}} and polynomial solutions of a system of linear equations”, Proc. ACM Symposium on Symbolic and Algebraic Computation, ACM (New York), 1986, pp. 222-224.

(3.12) T. Sasaki, “Simplification of algebraic expression by multiterm rewriting rules”, Proc. ACM Symposium on Symbolic and Algebraic Computation, ACM (New York), 1986, pp. 115-120.

(3.11) 佐々木建昭、元吉文男、 “国産数式処理システム GAL の開発”, 第２７回プログラミングシンポジウム報告集、情報処理学会、 1986, pp. 191-200.

(3.10) T. Sasaki, “Cramer-type formula for the polynomial solutions of coupled linear equations with polynomial coefficients”, Publ. RIMS. (Kyoto Univ.), Vol. 21, pp. 237-254 (1985).

(3.9) T. Sasaki and A. Furukawa, “Theory of multiple polynomial remainder sequence”, Publ. RIMS. (Kyoto Univ.), Vol. 20, pp. 367-399 (1984).

(3.8) A. Furukawa and T. Sasaki, “Multi-polynomial remainder sequence and its application to linear Diophantine equations”, Lect. Notes in Comp. Sci., Vol. 162, pp. 24-35 (1983).

(3.7) T. Sasaki and A. Furukawa, “Secondary polynomial remainder sequence and an extension of subresultant theory”, J. Inf. Proces., Vol. 7, pp. 175-184 (1984).

(3.6) T. Sasaki, S. Watanabe and A. Furukawa, “Design of a general computer algebra system”, Suuriken-Koukyuuroku (Kyoto Univ.), No. 486, pp. 154-165 (1983).

(3.5) T. Sasaki, “Constructing Bezout’s determinant from Sylvester’s determinant”, J. Inf. Proces., Vol. 6, pp. 163-166 (1983).

(3.4) T. Sasaki and H. Murao, “Efficient Gaussian elimination method for symbolic determinants and linear systems”, ACM Trans. Math. Software, Vol. 8, pp. 277-289 (1982).

(3.3) T. Sasaki, Y. Kanada and S. Watanabe, “Calculation of discriminants of high degree equations”, Tokyo J. Math., Vol. 4, pp. 493-500 (1981).

(3.2) T. Sasaki and Y. Kanada, “Parallelism in algebraic computation and parallel algorithms for symbolic linear systems”, Proc. ACM Symposium on Symbolic and Algebraic Computation, ACM (New York), 1981, pp. 160-167.

(3.1) T. Sasaki and Y. Kanada, “Practical rules for optimizing large multivariate polynomials”, Proc. Intn’l Meeting on Programming and Mathematical Methods for Solving Physical Problems, Dubna (USSR), 1977.

[4] 数値計算に関する論文

(4.6) T. Sasaki and Y. Kanada, “Practically fast multiple precision evaluation of log(x)”, J. Inf. Proces., Vol. 5, pp. 247-250 (1982).

(4.5) Y. Kanada and T. Sasaki, “LISP-based “big-float” system is not slow”, ACM SIGSAM Bulletin, Vol. 15, pp. 13-19 (1981).

(4.4) T. Sasaki, “An arbitrary precision real arithmetic package in REDUCE”, Lect. Notes in Comp. Sci., Vol. 72, pp. 358-368 (1979).

(4.3) T. Sasaki, “Multidimensional systematic sampling”, J. Inf. Proces., Vol. 4, pp. 79-88 (1981).

(4.2) T. Sasaki, “Multidimensional Monte Carlo integration based on factorized approximation functions”, SIAM J. Numer. Anal., Vol. 15, pp. 938-952 (1978).

(4.1) T. Sasaki, “Multidimensional integration techniques for ill-behaved functions”, Proc. Fourth Intn’l Colloquium on Advanced Computing Methods in Theoretical Physics, Marseille (France), 1977, pp. 230-247.

[5] 高エネルギー物理学に関する論文

(5.12) T. Sasaki, T. Kaneko, T. Ishikawa, K. Kato, S. Kawabata and Y. Shimizu, “An algorithm for calculating traces of gamma matrices in a compact form”, New Computing Techniques in Physics Research II (Ed. D. Perret-Gallix), World Scientific (Singapore), 1992, pp. 739-744.

(5.11) T. Sasaki, “Japanese formula manipulation system GAL”, Computing in High Energy Physics ’91, (Ed. Y. Watase and F. Abe), Universal Academy Press (Tokyo), 1991, pp. 383-389.

(5.10) Y. Oyanagi, T. Kaneko, T. Sasaki, S. Kawabata and Y. Shimizu, “How to calculate one-loop diagrams”, Perspectives of Particle Physics (Ed. R. Sasaki, et al.), World Scientific (Singapore), 1988.

(5.9) T. Sasaki, “Automatic generation of Feynman graphs in QED”, J. Comput. Phys., Vol. 22, pp. 189-214 (1976).

(4.8) N. Murai and T. Sasaki, “Quantum-number flow in multiparticle production”, Phys. Rev., Vol. D11, pp. 1083-1093 (1975).

(5.7) N. Murai and T. Sasaki, “New expression for particle distribution due to energy momentum conservation in longitudinal phase space”, Progr. Theor. Phys., Vol. 54, pp. 154-165 (1975).

(5.6) T. Sasaki and N. Murai, “Effect of resonance production in inclusive reactions and vector meson dominance model”, Phys. Rev., Vol. D9, pp. 3070-3082 (1974).

(5.5) T. Sasaki and N. Murai, “Slow onset of the scaling limit and resonance dominance”, Nuovo Cim. Letters, Vol. 7, pp. 129-132 (1973).

(5.4) N. Murai and T. Sasaki, “Inclusive reaction γP → π^{−} + X^{++} mediated by vector meson production”, Progr. Theor. Phys., Vol. 50, pp. 610-625 (1973).

(5.3) T. Sasaki, “A model of multiple production based on the intuitive picture of hadrons. II”, Progr. Theor. Phys., Vol. 49, pp. 1362-1363 (1973).

(5.2) T. Sasaki, “A model of multiple production based on intuitive picture of hadrons”, Progr. Theor. Phys., Vol. 48, pp. 198-209 (1972).

(5.1) T. Sasaki, “A possible existence of the second kind of weak interaction”, Progr. Theor. Phys., Vol. 45, pp. 854-865 (1971).

[6] 電子ビーム工学に関する論文

(6.9) T. Sasaki and M. Sasaki, “A theory of many-body Coulomb force I … Isotropic assembly”, Paper presented at Intn’l Conf. on Ion, Electon and Photon Beams, Los Angeles, May 1988.

(6.8) T. Sasaki, “Upper bound of the beam energy broadening in acceleration region”, J. Vac. Sci. Technol., Vol. B4, pp. 135-139 (1986).

(6.7) T. Sasaki, “Local space-charge forces and broadening of energy distribution in electron beams”, J. Vac. Sci. Technol., Vol. A2, pp. 1352-1360 (1984).

(6.6) T. Sasaki, “Trajectory displacement due to local space-charge force”, J. Vac. Sci. Technol., Vol. B1, pp. 1307-1311 (1983).

(6.5) T. Sasaki, “Study of space charge effect by computer”, J. Vac. Sci. Technol., Vol. 21, pp. 695-706 (1982).

(6.4) T. Sasaki, “A multibeam scheme for electron-beam lithography”, J. Vac. Sci. Technol., Vol. 19, pp. 963-965 (1981).

(6.3) T. Sasaki, “Effect of electron-electron interaction in computer simulation”, Proc. Ninth Intn’l Conf. on Electron and Ion Beam Science and Technology, St. Louis (USA), 1980, pp. 73-83.

(6.2) M. Idesawa, E. Goto, T. Soma and T. Sasaki, “Triple slit scheme for high current variable-shaped beam forming”, J. Vac. Sci. Technol., Vol. 16, pp. 1672-1675 (1979).

(6.1) T. Sasaki, “A computer study of electron-electron interaction in high density electron beams”, Proc. VLSI Conference, CALTEC (USA), 1979, pp. 125-137.

[7] コンピュータアートに関する論文

(7.5) M. K. Sasaki and T. Sasaki, “Techniques in computer art”, Visual Computer, Vol. 2, pp. 106-111 (1986).

(7.4) M. K. Sasaki and T. Sasaki, “Towards an intelligent computer art system”, Picture Engineering (Springer Series on Information Sciences), Springer-Verlag (Tokyo), 1982, pp. 286-302.

(7.3) M. K. Sasaki and T. Sasaki, “Generating mathematical patterns in computer art - Part 1”, Computers and People, Nov-Dec. issue, pp. 7-11 (1981); “ ⋯ - Part 2”, Jan-Feb. issue, pp. 17-18 (1982).

(7.2) M. K. Sasaki and T. Sasaki, “Computer art system ART-3”, Computer Graphics and Art, Aug. issue, pp. 4-11 (1978).

(7.1) M. K. Sasaki and T. Sasaki, “Computers and beauty”, Creative Computing, Vol. 2, pp. 48-51 (1976).

[8] その他のテーマに関する論文

(8.1) Y. Arimoto, Y. Homma, T. Sasaki, K. Tomono and M. Kawai, “Analytical studies on biological activity of copper compounds”, BCPC Monograph, No. 31, pp. 211-214 (1985).

[9] 解説・総説

(9.12) 佐々木建昭・加古富志雄、 “ 『近似代数』とは？”、数理科学、サイエンス社、 1998 年 11 月号、 pp. 8-20 (1998).

(9.11) 佐々木建昭、 “数式処理の最新事情”、数理科学、サイエンス社、 1998 年 11 月号、 pp. 5-7 (1998).

(9.10) T. Sasaki, “Computer Algebra: where we are ? where we go ?”, Proc. ATCM’95, pp. 42-51.

(9.9) 佐々木建昭、 “数式処理：研究道具として、研究テーマとして”、応用数理（日本応用数理学会誌）、第 1 巻、 pp. 35-52 (1991).

(9.8) 野田松太郎・佐々木建昭・鈴木正幸、 “数値計算と数式処理の融合による精度保証”、情報処理（日本情報処理学会誌）、第 31 巻、 pp. 1204-1211 (1990).

(9.7) 佐々木建昭、 “近似的代数計算法”、数理解析研究所講究録、第 676 巻、 pp. 307-319 (1988).

(9.6) 佐々木建昭・古川昭夫、 “コンピュータ環論”、情報処理（日本情報処理学会誌）、第 27 巻、 pp. 404-413 (1986).

(9.5) 元吉文男・佐々木建昭、 “数式処理の歴史と将来の展望”、情報処理（日本情報処理学会誌）、第 27 巻、 pp. 362-370 (1986).

(9.4) 佐々木建昭、 “多重多項式剰余列の理論と応用”、コンピュータ・ソフトウェア（日本ソフトウェア科学会誌）、第 2 巻、 pp. 29-43 (1985).

(9.3) 佐々木建昭、 “計算機による数式処理”、システムと制御（日本自動制御協会誌）、第 26 巻、 pp. 201-208 (1982).

(9.2) 佐々木建昭、 “計算機による代数的計算”、数学セミナ、日本評論社、 1981 年 7 月号、 pp. 23-29 (1981).

(9.1) 佐々木建昭・後藤英一、 “計算機による数式処理と物理学への応用”、物理（日本物理学会誌）、第 35 巻、 pp. 99-107 (1980).